OpenStudy (anonymous):

Solve: −2(x − 8) = −10

OpenStudy (anonymous):

13

OpenStudy (anonymous):

-2(x-8)=-10 -2x+16=-10 -2x=-26 x=13

OpenStudy (anonymous):

divide by \(-2\) get \[x-8=5\] add 8 get \[x=13\]